sample 14 similar cars and find a mean price of $19,850 with a standard deviation of $1184 12. Find the test statistic (t.s.) for your sample. (the formula is: t.) a) 2.24 b) -2.24 c 16 d) 16 e) none of these Find the critical t-score (t ) for the two tailed test with α-05. a) ±1.96 13. b) ±2.16 c) ±1.77 d) ±1.76 e) none of these

14. Find the critical t-score (te) for the two tailed test with a .01. a) t2.58 b) 3.01 c) 2.62 d) 2.65 e) none of these Therefore, the conclusion for this hypothesis test is: a) reject Ho at α c) reject Ho at α 15. .05 and FTR Ho at α-01 .01 and FTR Ho at α .05 b) reject Ho at α-.05 and α-01 d) FTR Ho at α .05 and u-01 e) none of these of feit nrangred using the current process show discoloration
Answer #1

Given that, sample size ( n ) = 14

sample mean (ar x) = $19850

sample standard deviation ( s ) = $1184

The null and alternative hypotheses are,

Ho : μ $20500

Ha : μ $20500

Q.12) Test statistic is,

19850- 20500 1184 V14 2.05

t.s. = -2.05

Answer: e) none of these

Q.13) t-scores ( tc) for the two tailed test at a-0.05 with df = 14 -1= 13 is, tc = +/- 2.16

Answer: b) +/- 2.16

Q.14) t-scores ( tc ) for the two tailed test at a-0.05 with df = 14 - 1 = 13 are tc = +/- 3.01

Answer: b) +/- 3.01

Q.15)

t.s. = -2.05 > -2.16

So, we fail to reject the null hypothesis ( H0).

t.s. = -2.05 > -3.01

So, we fail to reject the null hypothesis ( H0).

Answer: d) FTR H0 at a-0.05 and α 0.01